OpenStudy (anonymous):

Do the graphs of the equations given represent parallel lines? Both explain and graph your answer. y = x + 7 x - y = -2

OpenStudy (cwrw238):

write them both in the form y = mx + c where m = slope of the line y = x + 7 AND y = x + 2 are the slopes of these 2 the same?

OpenStudy (cwrw238):

graphs with the same slope are of course parallel

OpenStudy (cwrw238):

no compare the above with y = = mx + c y = mx + c y = 1x + 7 y = 1x + 2 what is m in each case? Compare them term by term

OpenStudy (cwrw238):

m = ?

OpenStudy (anonymous):

ok thank you :) I get it now

OpenStudy (cwrw238):

good - the slope is 1 in both cases so the lines are parallel the values of c denote the intersection of the lines on the y axis ( when x = 0)

OpenStudy (cwrw238):